# GASE: Generatively Augmented Sentence Encoding

### v1

@@ 4 Discussion
Overall, lower non-augmented performance was associated with larger performance improvements through generative augmentation as visualized in Figure 3. We believe this relationship holds because weaker embedding models are inherently less capable of modelling diverse semantics and, therefore, benefit more from the semantic diversity introduced by generative models. Moreover, embedding models with lower baseline performance also gained more from the ensemble effect of using $k=2$ models for augmentation compared to a single generative model. The effectiveness of keyword extraction for GloVe may be explained by the model’s limited capability to handle complex semantics as its sentence embeddings are derived by calculating an unweighted average of word embeddings. Therefore, the reduction to keywords might help the model to reduce noise.
We believe the lower scores on STS22 can be explained by its higher word count. Unexpectedly, generating summaries did not significantly reduce the mean word count on datasets other than STS22 (and BIOSSES, to some extent). Since most MTEB STS datasets include shorter sentences, sentence encoders might be optimized for shorter sentence lengths. Our results demonstrate that the best augmentation strategy for a given embedding model is not identical for all datasets.
@@ 5 Conclusion
We propose Generatively Augmented Sentence Encoding (GASE), a novel approach that augments sentence embeddings by paraphrasing, summarising, or extracting keywords with generative models at inference time. On the MTEB STS task, GASE leads to substantial improvements for embedding models with lower baseline performance and incremental improvements for SOTA embedding models. GASE is widely applicable as it only requires access to the outputs of the underlying embedding and generative models.
@@ 6 Limitations
  2. 2.
    Our investigation was confined to the evaluation of GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini for generative augmentation. Alternative LLMs, including the Claude and Gemini model families, may potentially offer more substantial improvements.
@@
    Furthermore, our experimental design was restricted to $k\in\{0,1,2\}$. While values of $k>2$ were not explored, we posit that such configurations may be impractical for real-world applications due to prohibitive computational costs.
@@
  5. 5.
    Our experiments are limited to the English language and the effectiveness of GASE for other languages requires further investigation. Since embedding models with weaker baseline performance showed larger improvements through generative augmentation, we hypothesise that GASE can provide larger performance improvements for languages for which the performance of embedding models is below English.
@@ Acknowledgments
Appendix A Prompts
@@
Appendix B Additional Experimental Results
@@
bert-large-cased - 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
@@
embed-english-light-v3.0 - 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
@@
voyage-lite-02-instruct - 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
@@
voyage-large-2-instruct - 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
@@
mxbai-embed-large-v1 - 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
@@
Table 4: Scores per dataset with and without generative augmentation using GPT-3.5 Turbo (Spearman’s rank correlation in %, bold: highest scores).
@@
GPT-4o mini 60.04 55.09 60.12 72.65 62.45 71.60 66.49 63.39 55.48 60.36 32.76
gpt-3.5-turbo, reka-flash 61.97 59.46 60.06 72.01 62.68 72.46 67.62 65.43 61.46 61.51 36.95
bert-large-cased - 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
gpt-3.5-turbo 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
@@
GPT-4o mini 65.01 65.37 54.27 68.60 58.18 72.26 71.12 73.54 55.72 70.50 60.54
gpt-3.5-turbo, reka-flash 68.01 69.39 56.86 70.04 62.30 74.61 73.86 74.52 60.49 71.32 66.72
all-MiniLM-L6-v2 - 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
@@
GPT-4o mini 79.81 81.97 74.03 83.31 76.70 86.03 79.19 87.69 67.01 80.08 82.11
gpt-3.5-turbo, reka-flash 80.64 82.30 74.59 83.72 78.11 86.20 79.95 87.23 69.23 80.12 84.97
all-mpnet-base-v2 - 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
@@
GPT-4o mini 81.07 83.75 74.19 85.19 79.28 86.45 80.47 90.89 67.81 81.23 81.43
gpt-3.5-turbo, reka-flash 81.72 84.43 75.46 86.12 80.63 87.00 81.21 89.37 69.33 81.55 82.13
embed-english-light-v3.0 - 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
@@
GPT-4o mini 79.90 84.14 75.76 79.65 77.89 85.12 79.83 87.90 67.91 78.68 82.10
gpt-3.5-turbo, reka-flash 80.54 84.48 76.45 80.61 79.65 85.54 81.21 87.60 69.23 79.06 81.55
embed-english-v3.0 - 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
@@
GPT-4o mini 81.82 86.52 76.80 82.57 80.11 87.77 83.72 89.42 68.36 77.90 85.05
gpt-3.5-turbo, reka-flash 82.53 86.68 77.88 83.56 81.64 88.05 84.13 88.59 69.87 78.44 86.48
voyage-2 - 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
@@
GPT-4o mini 83.07 86.89 79.59 87.27 81.23 88.46 85.81 89.37 65.25 80.85 86.02
gpt-3.5-turbo, reka-flash 83.80 87.27 80.62 87.99 82.40 88.51 86.40 88.77 67.94 80.87 87.28
voyage-lite-02-instruct - 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
@@
GPT-4o mini 85.97 88.25 85.24 89.06 86.62 89.57 85.99 86.90 78.59 79.45 90.08
gpt-3.5-turbo, reka-flash 86.05 88.74 84.83 89.14 86.69 89.36 86.57 86.72 78.81 79.34 90.34
voyage-large-2 - 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
@@
GPT-4o mini 83.95 87.54 80.02 87.91 83.22 89.11 85.68 90.47 63.99 81.54 90.08
gpt-3.5-turbo, reka-flash 84.12 87.56 80.05 88.31 83.98 88.76 85.97 90.04 64.84 81.55 90.14
voyage-large-2-instruct - 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
@@
GPT-4o mini 84.78 88.83 78.71 88.43 86.45 91.17 85.89 89.14 67.09 83.19 88.87
gpt-3.5-turbo, reka-flash 84.86 89.16 79.11 88.56 86.67 90.97 85.86 89.67 67.56 83.00 88.08
mxbai-embed-large-v1 - 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13

### v2

@@ 4 Discussion
We believe that embedding models with lower baseline performance benefited more from the semantic diversity induced by augmentation due to their inherent lower capability to model diverse semantics. Similarly, they gained more from the ensemble effect using $k\geq 2$ generative models. Our extension to $11$ non-English languages and the PC and IR tasks did confirm the effectiveness of GASE and also showed a negative association between baseline model performance and improvements through augmentation.
Figure 3: Avg. STS scores for differently weighted averages between embeddings of paraphrases generated with GPT-3.5-Turbo and original texts ($k=1$).
The effectiveness of keyword extraction for GloVe, BERT, and Llama-3 may be explained by the model’s limited capability to handle complex semantics. Therefore, the reduction to keywords might help the model to reduce noise.
Longer texts likely caused summary augmentation to outperform paraphrase augmentation on STS22, as summarising significantly reduced the length. On other datasets, however, summaries were only slightly shorter, diminishing its augmentation value. Moreover, the lower Jaccard Similarity between paraphrases and the original texts indicates greater variation, which in turn makes embeddings more effective. Consistent with this, summaries outperform paraphrases on STS22 (Table 1) - the only dataset where summaries exhibit a lower Jaccard Similarity than paraphrases (Table 4). Our results show pooling original and synthetic embeddings outperforms individual use (Figure 3). Across datasets, tasks, and languages, GASE particularly benefits embedding models with lower baseline performance and shorter texts, while summary augmentation is more effective for longer texts (Table 1). Moreover, GASE requires a certain size of the generative model to be effective (Table 8). For IR, we found that augmenting both the queries and the corpus is necessary (Table 9) which is particularly costly due to document length.
Since GenEOL used different, mostly weaker, embedding models, our results cannot be directly compared with their results.
@@ 5 Conclusion
We propose GASE, which augments sentence encoders at inference with generative models for paraphrasing, summarising, or keyword extraction. On MTEB STS, GASE significantly improves lower-performing embedding models and incrementally enhances SOTA models. GASE applies broadly, needing only embedding and generative model outputs. Extensions to $11$ non-English languages and Pair Classification and Information Retrieval as additional tasks, demonstrated consistent findings.
@@ 6 Limitations
    Our approach significantly increases the computational cost for sentence encoding since each generative augmentation requires an additional inference by the embedding model plus inference using a generative LLM. For a basic estimate of the runtime increase based on STS17 see Table 20. The results highlight GASE’s particular suitability for smaller embedding models, where it yields significant performance gains with minimal encoder runtime.
  2. 2.
@@
    Our experiments demonstrate that the effectiveness of GASE is limited to embedding models with lower baseline performance while encoders with higher baseline performance exhibit minimal to no improvement. Further investigation is needed to understand the underlying reasons for this observation.
@@
    Our investigation was confined to evaluating GPT-3.5 Turbo, Reka-Flash, GPT-4o mini, and to some limited extent Qwen2.5 for generative augmentation. Alternative LLMs, including the Claude and Gemini model families, may lead to different results.
  5. 5.
    Furthermore, our experimental design was restricted to $k\in\{0,1,2,3\}$. While values of $k>3$ were not explored, we posit that such configurations may be impractical for real-world applications due to prohibitive computational costs.
  6. 6.
    Our experiments for generative model ablation are limited to models with a parameter count $\in\{0.5b,1.5b,3b,7b,14b\}$. Extending the analysis to larger generative models would yield insights into whether the observed pattern holds for models with more than 1414b parameters.
  7. 7.
@@ Acknowledgments
This research was partially supported by the Horizon Europe project GenDAI (Grant Agreement ID: 101182801) and by the ADAPT Research Centre at Munster Technological University. ADAPT is funded by Taighde Éireann – Research Ireland through the Research Centres Programme and co-funded under the European Regional Development Fund (ERDF) via Grant 13/RC/2106_P2.
We would also like to thank the anonymous reviewers for their valuable feedback and constructive suggestions, which have helped to improve the quality and clarity of this work.
Appendix A Comparison of GASE and GenEOL
A comparison of our approach, GASE, and GenEOL can be found in Table 11.
GASE (ours) GenEOL
Transformations Paraphrasing
Summarising
Extracting keywords
Paraphrasing
Consice paraphrasing
Changing the sentence structure
Entailment
Each one followed by summarisation.
Number of transformations $k\in\{0,1,2,3\}$ $m\in\{0,2,4,8,16,24,32\}$
Pooling Mean pooling over variations generated using $k$ different generative LLMs Mean pooling over variations generated using $m$ different transformations
Embeding models English-language models:
glove.840B.300d
bert-large-cased
all-MiniLM-L6-v2
all-mpnet-base-v2 embed-english-light-v3.0
embed-english-v3.0
voyage-2
voyage-lite-02-instruct
voyage-large-2 voyage-large-2-instruct
mxbai-embed-large-v1
llama-3-8b
Multilingual models:
paraphrase-multilingual-
mpnet-base-v2
multilingual-e5-large-
instruct
voyage-multilingual-2
mistral0.1-7B
llama-2-7B
llama-3-8B
bert-large
Generative models gpt-3.5-turbo-0125
reka-flash-20240226
gpt-4o-mini-2024-07-18
qwen2.5-0.5b
qwen2.5-1.5b
qwen2.5-3b
qwen2.5-7b
qwen2.5-14b
qwen2.5-32b
gpt-3.5-turbo-0125
mistral0.1-I-7B
STS datasets Complete English MTEB for STS:
STSB
STS12
STS13
STS14
STS15
STS16
STS17
STS22
SICK-R
BIOSSES
Multilingual datasets:
STS17
STS22
Subset of English MTEB for STS:
STSB
STS12
STS13
STS14
STS15
STS16
SICK-R
Other tasks Text Pair Classification
Retrieval
Text Pair Classification
Classification
Clustering
Reranking
Table 11: Comparison of GASE and GenEOL.
Based on $7$ STS datasets, Thirukovalluru and Dhingra (2025) report a Spearman’s rank correlation for the weakest baseline model of $47.06$ and $80.37$ for their best model using $32$ generative augmentations. Using the same $7$ datasets, we find that without any augmentation their best performing model is outperformed by $7$ of the $12$ embedding models evaluated in this work (e.g. by all-mpnet-base-v2; see Table 15 and Table 16).
Appendix B Random keyword extraction
Random keyword extraction randomly selects $16.1/58.2\approx 0.28$ of the words from a text sequence. This share is equal to the average share of the number of LLM-extracted keywords and the total word count of a sentence (see Table 3). As this results in a low number of keywords for short sentences, the minimal number of randomly selected keywords was set to $3$. All punctuation has been removed before extracting the keywords.
Appendix C Prompts
@@
Model-specific post-processing is performed to remove meta-text not used for augmentation.
Appendix D Embeddings models
Embedding models for the English language:
  • glove.840B.300d (Pennington et al., 2014)
  • bert-large-cased (Devlin et al., 2019)
  • all-MiniLM-L6-v2 Sentence Transformers (2024)
  • all-mpnet-base-v2 Sentence Transformers (2024)
  • embed-english-light-v3.0 Cohere (2024)
  • embed-english-v3.0 Cohere (2024)
  • voyage-2 (Voyage AI, 2024)
  • voyage-lite-02-instruct (Voyage AI, 2024)
  • voyage-large-2 (Voyage AI, 2024)
  • voyage-large-2-instruct (Voyage AI, 2024)
  • mxbai-embed-large-v1 Mixedbread (2024)
  • llama-3-8b Meta (2024)
For llama-3-8b we used the following KEEOL Zhang et al. (2024) prompt:
The essence of a sentence is often captured by its main subjects and actions, while descriptive terms provide additional but less central details. With this in mind, this sentence: “input_text” means in one word:
Following Zhang et al. (2024), we used the penultimate layer of llama-3-8b to extract the embeddings.
Multilingual embedding models:
  • paraphrase-multilingual-mpnet-
    base-v2
     Reimers and Gurevych (2020)
  • intfloat/multilingual-e5-
    large-instruct
     Wang et al. (2024)
  • voyage-multilingual-2 Voyage AI (2024)
Appendix E Datasets
The following datasets from the MTEB STS leaderboard (Hugging Face, 2024) were used:
  • STSB (Cer et al., 2017)
  • STS12 (Agirre et al., 2012)
  • STS13 (Agirre et al., 2013)
  • STS14 (Agirre et al., 2014)
  • STS15 (Agirre et al., 2015)
  • STS16 (Agirre et al., 2016)
  • STS17 (Cer et al., 2017)
  • STS22 (Chen et al., 2022)
  • SICK-R Marelli et al. (2014)
  • BIOSSES (Soğancıoğlu et al., 2017)
The original MTEB paper (Muennighoff et al., 2023) also included the STS11 dataset in Figure 1 of their paper where the authors specify the MTEB datasets, but excluded it in the evaluations. Similarly, the MTEB Leaderboard (Hugging Face, 2024) does not include STS11 either. To make our results comparable to other models evaluated on MTEB we therefore excluded STS11 from our evaluations. All data has been obtained from the MTEB repository on the Hugging Face website. [^13] https://huggingface.co/mteb/datasets
The following datasets were used for our extensions:
  • Twitter SemEval 2015 (Xu et al., 2015)
  • Twitter URL-Corpus (Lan et al., 2017)
  • Sprint duplicate questions (Shah et al., 2018)
  • NFCorpus (Boteva et al., 2016)
Appendix F Generative Model Versions, Hyperparameters, and Hardware Configuration
The following versions have been used for generative augmentation:
  • GPT-3.5 Turbo: gpt-3.5-turbo-0125
  • Reka-Flash: reka-flash-20240226
  • GPT-4o mini: gpt-4o-mini-2024-07-18
  • Qwen2.5 (all models running via Ollama with quantization Q4_K_M):
    • qwen2.5:0.5b
    • qwen2.5:1.5b
    • qwen2.5:3b
    • qwen2.5:7b
    • qwen2.5:14b
GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini have been accessed through the respective APIs with temperature set to 0 and top_p to $1$ to make the experiments as reproducible as possible. Where applicable, a random seed of $1337$ was used. Other hyperparameters were used with default values.
All Qwen2.5 model runs were conducted with Ollama on a local computer with Q4_K_M quantisation and the following specification:
  • GeForce RTX 3090 MSI Gaming X Trio 24GB GPU
  • AMD Ryzen 9 9950X CPU
  • DDR5-6400 64GB RAM
Appendix G Additional Experimental Results
Embedding model Pooling STS Score
glove.840B.300d
mean 59.98
max 59.02
bert-large-cased
mean 66.29
max 65.16
all-MiniLM-L6-v2
mean 80.32
max 79.88
all-mpnet-base-v2
mean 81.39
max 81.03
embed-english-
light-v3.0
mean 80.30
max 79.74
embed-english-v3.0
mean 82.10
max 81.62
voyage-2
mean 83.38
max 82.69
voyage-lite-02-
instruct
mean 85.99
max 85.53
voyage-large-2
mean 83.92
max 83.48
voyage-large-2-
instruct
mean 84.80
max 84.41
mxbai-embed-
large-v1
mean 85.01
max 84.69
llama-3-8b
mean 75.28
max 75.07
Table 12: Average STS scores across tasks for different pooling methods in % (bold: best per embedding model).
@@ Appendix G Additional Experimental Results
bert-large-cased none 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
@@
embed-english-light-v3.0 none 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
@@
voyage-lite-02-instruct none 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
@@
voyage-large-2-instruct none 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
@@
mxbai-embed-large-v1 none 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
@@
llama-3-8b none 71.36 79.79 62.81 79.08 69.72 78.49 80.88 84.12 33.39 76.27 69.04
paraphrasing 75.28 81.37 67.03 81.29 72.88 80.91 82.60 83.04 50.55 78.25 74.84
summarising 76.74 80.39 65.02 81.18 74.51 82.07 81.00 84.48 59.44 79.36 79.92
extracting keywords 74.95 78.57 67.39 82.68 74.74 81.67 80.31 85.71 51.79 71.05 75.60
Table 13: Scores per dataset with and without generative augmentation using GPT-3.5 Turbo in % (bold: highest scores).
Embedding
model
Augment-
tation
Stop-
word
removal
Average STSB STS12 STS13 STS14 STS15 STS16 STS17 STS22 SICK-R BIOSSES
glove.840B.
300d
none no 57.86 50.73 57.50 70.98 60.69 70.85 63.85 62.05 54.08 55.42 32.45
yes 60.07 57.13 57.43 70.91 60.68 70.86 63.84 75.83 54.66 54.97 34.39
para-
phrasing
no 59.98 56.78 57.33 71.17 61.16 71.52 67.66 62.28 57.78 58.92 35.21
yes 65.43 67.05 59.98 72.72 63.75 73.70 69.98 75.01 61.01 61.94 49.17
bert-large-cased none no 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
yes 53.08 50.11 41.37 53.10 50.11 63.68 59.66 63.73 49.69 48.75 50.54
para-
phrasing
no 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.79 64.56
yes 60.25 61.42 50.45 58.43 55.20 69.75 67.19 67.98 57.01 58.25 56.82
all-MiniLM-
L6-v2
none no 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
yes 75.36 78.80 68.29 77.62 73.05 84.35 74.92 87.05 65.74 65.58 78.24
para-
phrasing
no 80.32 82.33 73.33 83.33 77.85 86.23 79.88 87.02 68.47 79.63 85.07
yes 77.78 80.69 71.61 81.15 75.92 84.86 77.51 85.89 66.36 70.61 83.21
all-mpnet-base-
v2
none no 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
yes 75.51 78.32 65.71 79.49 73.99 83.98 73.28 88.98 65.82 67.01 78.47
para-
phrasing
no 81.39 84.10 74.04 85.52 80.52 86.82 81.15 89.08 69.01 81.25 82.39
yes 78.20 80.87 70.76 83.09 77.53 85.00 76.72 87.05 66.28 71.99 82.70
embed-english-
light-v3.0
none no 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
yes 74.36 78.16 68.36 76.23 73.84 81.67 71.69 88.18 66.87 64.32 74.34
para-
phrasing
no 80.30 84.65 74.65 80.73 79.41 85.45 81.14 87.45 68.42 78.75 82.36
yes 77.58 81.43 73.12 79.15 76.81 83.77 76.89 87.55 67.81 70.40 78.87
embed-english-
v3.0
none no 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
yes 76.70 79.52 67.15 80.84 75.37 83.76 74.40 88.51 66.99 66.56 83.95
para-
phrasing
no 82.10 86.69 76.37 83.71 81.29 87.96 83.95 88.47 68.63 78.21 85.75
yes 79.26 83.00 72.79 83.09 78.72 85.21 79.03 87.36 68.03 71.45 83.96
voyage-2 none no 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
yes 78.11 81.52 71.71 82.73 77.23 85.59 79.12 87.91 66.40 67.22 81.64
para-
phrasing
no 83.38 87.17 79.23 87.13 81.67 88.57 86.27 89.07 66.80 80.33 87.56
yes 80.42 84.16 76.32 84.89 79.29 86.33 82.93 87.58 67.55 71.42 83.69
voyage-lite-02-
instruct
none no 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
yes 81.35 83.24 79.33 86.13 82.37 87.95 80.01 85.84 74.72 66.73 87.14
para-
phrasing
no 85.99 88.81 84.47 89.05 86.68 89.52 86.53 86.79 79.08 79.09 89.89
yes 82.88 85.57 81.64 87.14 83.29 87.58 83.46 85.18 76.02 70.83 88.11
voyage-large-2 none no 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
yes 79.39 82.25 72.85 82.94 78.99 86.54 79.68 89.63 65.38 67.42 88.25
para-
phrasing
no 83.92 87.62 79.18 87.66 83.60 88.91 86.00 90.22 64.69 81.00 90.29
yes 81.19 84.57 76.72 84.73 80.84 87.17 82.95 88.78 65.43 71.64 89.04
voyage-large-2-
instruct
none no 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
yes 81.35 83.24 79.33 86.13 82.37 87.95 80.01 85.84 74.72 66.73 87.14
para-
phrasing
no 84.80 89.31 77.61 88.40 86.62 91.17 86.20 89.58 67.72 82.83 88.58
yes 81.88 85.70 74.85 87.76 83.91 88.69 82.40 88.41 67.66 71.72 87.73
mxbai-embed-
large-v1
none no 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13
yes 79.54 83.54 72.88 85.73 81.33 86.72 78.69 86.36 67.41 68.76 83.99
para-
phrasing
no 85.01 89.09 80.83 89.43 85.88 89.46 86.94 88.75 70.03 82.67 87.02
yes 81.36 86.15 78.11 87.20 83.03 86.93 78.69 86.40 68.47 72.70 85.97
llama-3-8b
none no 71.36 79.79 62.81 79.08 69.72 78.49 80.88 84.12 33.39 76.27 69.04
yes 69.50 73.48 59.66 77.37 69.66 77.17 72.94 83.00 50.94 62.15 68.58
para-
phrasing
no 75.28 81.37 67.03 81.29 72.88 80.91 82.60 83.04 50.55 78.25 74.84
yes 74.57 78.29 66.98 80.90 73.61 79.63 78.76 83.86 61.88 68.92 72.90
Table 14: Scores per dataset with and without stopword removal using either no generative augmentation or paraphrase augmentation with GPT-3.5 Turbo in %.
@@
gpt-4o-mini 60.04 55.09 60.12 72.65 62.45 71.60 66.49 63.39 55.48 60.36 32.76
gpt-3.5-turbo
+ reka-flash
61.97 59.46 60.06 72.01 62.68 72.46 67.62 65.43 61.46 61.51 36.95
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
62.18 59.99 61.79 72.88 64.04 72.03 67.38 64.75 60.76 61.79 36.42
bert-large-cased none 60.78 59.18 45.71 63.94 54.42 68.98 65.46 71.30 55.48 64.16 59.14
gpt-3.5-turbo 66.29 67.97 53.62 68.56 59.73 74.01 71.97 74.14 58.57 69.80 64.56
@@
gpt-4o-mini 65.01 65.37 54.27 68.60 58.18 72.26 71.12 73.54 55.72 70.50 60.54
gpt-3.5-turbo
+ reka-flash
68.01 69.39 56.86 70.04 62.30 74.61 73.86 74.52 60.49 71.32 66.72
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
68.41 69.40 59.01 71.25 62.60 74.13 73.86 74.99 60.35 71.74 66.73
all-MiniLM-L6-v2 none 78.91 82.03 72.37 80.60 75.59 85.39 78.99 87.59 67.26 77.58 81.64
@@
gpt-4o-mini 79.81 81.97 74.03 83.31 76.70 86.03 79.19 87.69 67.01 80.08 82.11
gpt-3.5-turbo
+ reka-flash
80.64 82.30 74.59 83.72 78.11 86.20 79.95 87.23 69.23 80.12 84.97
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
80.76 82.31 75.16 84.29 78.22 86.18 79.98 87.31 68.84 80.34 85.02
all-mpnet-base-v2 none 80.28 83.42 72.63 83.48 78.00 85.66 80.03 90.60 68.00 80.59 80.43
@@
gpt-4o-mini 81.07 83.75 74.19 85.19 79.28 86.45 80.47 90.89 67.81 81.23 81.43
gpt-3.5-turbo
+ reka-flash
81.72 84.43 75.46 86.12 80.63 87.00 81.21 89.37 69.33 81.55 82.13
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
81.85 84.42 75.96 86.37 80.72 86.92 81.33 89.68 69.05 81.50 82.51
embed-english-light-v3.0 none 78.75 83.52 72.81 77.69 76.91 83.51 78.49 88.49 67.88 77.98 80.20
@@
gpt-4o-mini 79.90 84.14 75.76 79.65 77.89 85.12 79.83 87.90 67.91 78.68 82.10
gpt-3.5-turbo
+ reka-flash
80.54 84.48 76.45 80.61 79.65 85.54 81.21 87.60 69.23 79.06 81.55
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
80.66 84.59 77.10 81.00 79.75 85.68 81.16 87.70 68.90 79.09 81.61
embed-english-v3.0 none 81.25 86.54 74.76 81.68 78.81 87.18 83.01 89.70 68.22 77.52 85.05
@@
gpt-4o-mini 81.82 86.52 76.80 82.57 80.11 87.77 83.72 89.42 68.36 77.90 85.05
gpt-3.5-turbo
+ reka-flash
82.53 86.68 77.88 83.56 81.64 88.05 84.13 88.59 69.87 78.44 86.48
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
82.08 86.69 78.49 83.60 81.65 87.97 84.10 88.76 69.67 78.35 86.36
Table 15: Scores per dataset with and without paraphrase augmentation using GPT-3.5 Turbo, Reka-Flash, and GPT-4o mini in % (bold: highest scores) (part 1).
voyage-2 none 82.50 87.06 77.68 86.25 80.11 88.38 85.72 89.78 65.26 78.95 85.83
@@
gpt-4o-mini 83.07 86.89 79.59 87.27 81.23 88.46 85.81 89.37 65.25 80.85 86.02
gpt-3.5-turbo
+ reka-flash
83.80 87.27 80.62 87.99 82.40 88.51 86.40 88.77 67.94 80.87 87.28
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
83.72 87.14 81.06 88.14 82.50 88.40 86.14 88.81 66.91 81.00 87.08
voyage-lite-02-instruct none 85.95 88.74 86.19 88.84 86.84 89.84 86.07 87.19 78.62 77.54 89.60
@@
gpt-4o-mini 85.97 88.25 85.24 89.06 86.62 89.57 85.99 86.90 78.59 79.45 90.08
gpt-3.5-turbo
+ reka-flash
86.05 88.74 84.83 89.14 86.69 89.36 86.57 86.72 78.81 79.34 90.34
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
86.01 88.48 84.95 89.26 86.74 89.25 86.32 86.75 78.31 79.50 90.55
voyage-large-2 none 83.60 87.84 78.66 86.98 82.25 88.94 85.71 91.29 63.97 79.83 90.49
@@
gpt-4o-mini 83.95 87.54 80.02 87.91 83.22 89.11 85.68 90.47 63.99 81.54 90.08
gpt-3.5-turbo
+ reka-flash
84.12 87.56 80.05 88.31 83.98 88.76 85.97 90.04 64.84 81.55 90.14
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
84.17 87.49 80.53 88.42 84.00 88.73 85.77 89.82 64.69 81.75 90.49
voyage-large-2-instruct none 84.61 89.21 76.15 88.49 86.50 91.13 85.68 90.05 66.59 83.17 89.09
@@
gpt-4o-mini 84.78 88.83 78.71 88.43 86.45 91.17 85.89 89.14 67.09 83.19 88.87
gpt-3.5-turbo
+ reka-flash
84.86 89.16 79.11 88.56 86.67 90.97 85.86 89.67 67.56 83.00 88.08
gpt-3.5-turbo
+ reka-flash
+ gpt-4o-mini
84.91 88.98 79.79 88.59 86.69 90.89 85.83 89.23 67.62 82.97 88.51
mxbai-embed-large-v1 none 84.63 89.29 79.07 89.80 85.22 89.34 86.77 89.21 68.73 82.78 86.13